In a survey of 9000 women, 6431 sy they change their nail polish once a week. Construct a 95% confidence interval for the population proportion of women who change their nail polish once a week

9000/6431=1.4

Answer by stanbon(75887)

In a survey of 9000 women, 6431 sy they change their nail polish once a week. Construct a 95% confidence interval for the population proportion of women who change their nail polish once a week
------
sample proportin = p-hat = 6431/9000 = 0.71
----
ME = 1.96*sqrt[0.71*0.29/9000] = 0.0094
------
95% CI: 0.71-0.0094 < p < 0.71+0.0094
==========================================
Cheers,
Stan H.
================